OpenStudy (anonymous):

In a triangle ABC, if a=18 , b=24, c=30 find sinA, sinB ,sinC.

OpenStudy (anonymous):

law o' cosines

OpenStudy (anonymous):

\[a^2=b^2+c^2-2bc\cos(A)\] \[\cos(A)=\frac{b^2+c^2-a^2}{2bc}\] \[A=\cos^{-1}\left(\frac{b^2+c^2-a^2}{2bc}\right)\]